if two dice are rolled one time find the probability a 3 on one die or on both dice


1
Expert's answer
2021-11-09T17:18:51-0500

When two dice are rolled then total sample space = 6*6 = 36

There are total 10 outcomes when obtaining 3 on one die in case of rolling two dice. Following are these expected outcomes.

[(3,1), (3,2), (3,4), (3,5), (3,6), (1,3), (2,3), (4,3), (5,3), (6,3)]

There is only one outcome when obtaining 3 on both dice.

Then, the probability a 3 on one die or on both dice is:

"P(A) = \frac{10}{36}+\frac{1}{36}=\frac{11}{36}=0.306"